OpenStudy (anonymous):

what times 7 x N =50 please some one just GIVE me the answer!!!

OpenStudy (anonymous):

7.14285714285714

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

just divide lol

OpenStudy (anonymous):

\(\huge\color{blue}{Welcome~to~OpenStudy!}\)

OpenStudy (anonymous):

OK :D

OpenStudy (anonymous):

\(\huge\color{blue}{The~answer~is~7.14285714.}\)

OpenStudy (anonymous):

i already said that

sammixboo (sammixboo):

\(7*n=50\) \(7÷7*n=50÷7\) \(n=50÷7\) What is \(50÷7?\)

sammixboo (sammixboo):

^ That is how you answer it ^

OpenStudy (anonymous):

they already got the answer lol

sammixboo (sammixboo):

I know I was just explaining to him/her HOW they can get the answer for future times